# TrackGS: Optimizing COLMAP-Free 3D Gaussian Splatting with Global Track Constraints

### v2

@@ Abstract
While 3D Gaussian Splatting (3DGS) has advanced ability on novel view synthesis, it still depends on accurate pre-computaed camera parameters, which are hard to obtain and prone to noise. Previous COLMAP-Free methods optimize camera poses using local constraints, but they often struggle in complex scenarios. To address this, we introduce TrackGS, which incorporates feature tracks to globally constrain multi-view geometry. We select the Gaussians associated with each track, which will be trained and rescaled to an infinitesimally small size to guarantee the spatial accuracy. We also propose minimizing both reprojection and backprojection errors for better geometric consistency. Moreover, by deriving the gradient of intrinsics, we unify camera parameter estimation with 3DGS training into a joint optimization framework, achieving SOTA performance on challenging datasets with severe camera movements. Code and datasets will be released.
Figure 1: Comparisons on novel view synthesis and camera poses. We propose a novel model of 3D Gaussian Splatting without any known camera parameters by leveraging global track information. Compared with state-of-the-art methods, we provide not only higher rendering quality in novel view synthesis, but also more accurate estimation of camera poses on benchmark datasets, including the challenging real-world indoor and outdoor or synthetic scenes with complicated camera movements (the right column).
1 Introduction
@@
Given a collection of images from a 3D scene along with the corresponding camera intrinsic and extrinsic parameters, 3D Gaussian Splatting (3DGS) [11] can effectively represent the scene with a series of 3D Gaussians, and generate high-quality images from novel viewpoints. Due to its efficiency in training and superior performance in testing, 3DGS has become popular for a variety of applications including reconstruction, editing, and AR/VR etc. However, the effectiveness of 3DGS training relies on accurately pre-determined camera poses (i.e., camera extrinsics) and camera focal lengths (i.e., camera intrinsics). These parameters are typically derived using COLMAP [24] in advance. This preprocessing step is not only time-consuming but also impacts the training performance of 3DGS, particularly when dealing with complex camera movements and scenes.
Recent COLMAP-Free approaches [32, 8, 1, 16, 5, 6, 9] have tried to address this by adding local constraints. However, it limits these methods to handling only simpler scenes. They typically assume that the input data are sequential and that the focal length is known. When faced with more complex scenarios with very complicated camera movements, these methods generally perform poorly.
To address this problem, we introduce global track information to globally constrain geometric consistency and innovatively integrate it into 3DGS. Specifically, we select the 3D Gaussians that correspond to each track, where the track Gaussians are dynamically adjusted in size during the training process, shrinking to an infinitesimally small size, and approaching spatial points that are distributed near the surface. Using our novel designed 2D and 3D track losses, the reprojection and backprojection errors are explicitly minimized. Moreover, the remaining 3D Gaussians continue to function as before, and all 3D Gaussians are aggregated and constrained by the loss function. Additionally, for the first time, we theoretically derive the focal length gradients to achieve full differentiability of the pipeline, eliminating the pre-computation of all the necessary camera parameters, including both the intrinsics and extrinsics. It enables to unify the learning process for both camera parameters and 3DGS.
@@
  • We are the first to propose integrating track information with 3DGS, using global geometric constraints to simultaneously optimize camera parameters and 3DGS. To achieve this, we introduce 2D and 3D track losses to constrain reprojection and backprojection errors.
  • We propose a joint optimization framework, where for the first time, we derive the gradient of the camera intrinsics. Without relying on any known camera parameters, we achieve full differentiability for the entire pipeline, seamlessly integrating camera parameters estimation, including both intrinsics and extrinsics, with 3DGS training.
@@
2 Related Work
2.1 Novel View Synthesis
@@
Novel view synthesis is a foundational task in the computer vision and graphics, which aims to generate unseen views of a scene from a given set of images. Numerous methods have been developed to address this problem by approaching it as 3D geometry-based rendering, such as using meshes [7, 21, 22], MPI [13, 25, 35], point clouds [33, 30], etc.
@@
Figure 2: Overview. Given a set of images, our method obtains both camera intrinsics and extrinsics, as well as a 3DGS model. During the initialization, we extract the global tracks, and initialize camera parameters and Gaussians from image correspondences and monodepth with reprojection loss. We determine Gaussian kernels with recovered 3D track points, and then jointly optimize the parameters $K$, $T_{cw}$, 3DGS through the proposed global track constraints (i.e., $L_{t2d}$, $L_{t3d}$, and $L_{scale}$) and original photometric losses (i.e., $L_{1}$ and $L_{D-SSIM}$).
@@
Recently, Neural Radiance Fields (NeRF) [17] provide a novel solution to this problem by representing scenes as implicit radiance fields using neural networks, achieving photo-realistic rendering quality. Although having some works in improving efficiency [18, 15], the time-consuming training and rendering still limit its practicality. Alternatively, 3D Gaussian Splatting (3DGS) [11] models the scene as explicit Gaussian kernels, with differentiable splatting for rendering. Its improved real-time rendering performance, lower storage and efficiency, quickly attract more attentions.
2.2 Optimizing Camera Poses in NeRFs and 3DGS
@@
Although NeRF and 3DGS can provide impressive scene representation, these methods all need accurate camera parameters (both intrinsic and extrinsic) as additional inputs, which are mostly obtained by COLMAP [24]. When the prior is inaccurate or unknown, accurately estimating camera parameters and scene representations becomes crucial.
@@
In earlier studies, scene training and camera pose estimation relied solely on photometric constraints. iNeRF [32] refines the camera poses using a pre-trained NeRF model. NeRFmm [29] introduces a joint optimization approach that simultaneously estimates camera poses and trains the NeRF model. BARF [14] and GARF [2] propose a new positional encoding strategy to address the gradient inconsistency issues in positional embedding, achieving promising results. However, these methods only yield satisfactory optimization when the initial pose is very close to the ground truth, as photometric constraints alone can only enhance camera estimation quality within a limited range. Subsequently, SC-NeRF [8] minimizes a projected ray distance loss based on correspondence between adjacent frames. NoPe-NeRF [1] utilizes monocular depth maps as geometric priors and defines undistorted depth loss and relative pose constraints.
@@
Regarding 3D Gaussian Splatting, CF-3DGS [6] utilizes mono-depth information to refine the optimization of local 3DGS for relative pose estimation and subsequently learns a global 3DGS in a sequential manner. InstantSplat [5] targets sparse view scenes, initially employing DUSt3R [27] to create a densely covered, pixel-aligned point set for initializing 3D Gaussian models, and then implements a parallel grid partitioning strategy to accelerate joint optimization. Jiang et al. [10] develops an incremental method for reconstructing camera poses and scenes, but it struggles with complex scenes and unordered images. HT-3DGS [9] interpolates frames for training and splits the scene into local clips, using a hierarchical strategy to build 3DGS model. It works well for simple scenes, but fails with dramatic motions due to unstable interpolation and low efficiency.
@@
where $c_{i}$ and $\alpha_{i}$ are the color and opacity of the Gaussians. Similarly, the depth of the scene preceived of a pixel is,
@@
4 Method
Overview. Given a set of images =$\mathcal{I}=\{I_{i}\}_{i=1}^{M}$, with unknown extrinsic matrix $T_{cw,i}$ at each view and unknown intrinsic matrix denoted by $K$, our method aims to build a 3D Gaussian Splatting (3DGS) model while simultaneously estimating both the extrinsic and intrinsic matrices, as shown in Fig. 2. To achieve this goal, our key approach is to leverage the global track constraint to explicitly capture and enforce multi-view geometric consistency, which serves as the foundation for accurately estimating both the 3DGS model and the camera parameters. Specifically, during initialization, we construct Maximum Spanning Tree (Sec. 4.1) based on 2D matched feature points and extract global tracks. Then we initialize both the camera parameters and subsequent 3D Gaussians with the estimated 3D track points. Building on this, we propose an effective joint optimization method with three loss terms: 2D track loss, 3D track loss, and scale loss. The 2D and 3D track losses are minimized to ensure multi-view geometric consistency. The scale loss constrains the track Gaussians remain aligned with the scene’s surface while preserving the expressive capability of the 3DGS model. We derive and implement the differentiable components of the camera parameters, including both the extrinsic and intrinsic matrices. This allows us to apply the chain rule, enabling seamless joint optimization of the 3DGS model and the camera parameters.
4.1 Initialization
Global Tracks. We begin by extracting 2D feature points $\{p_{i}\}$ from each image $I$ and computing feature matches across all images using off-the-shelf algorithms [4, 23]. To organize these matches into global tracks, we first construct a Maximum Spanning Tree (MST) using Kruskal’s algorithm, where the node represents each image and the weight of each edge is determined by the number of feature matching pairs between two images. By traversing MST and the feature points, we use the Union-Find algorithm to extract global track $\mathcal{P}$ and remove short tracks for robustness.
@@ 4.1 Initialization
For each edge $(i,j)$ in the MST, we leverage the off-the-shelf monocular depth maps of images (i.e., DPT [19]), and convert correspondences $p_{i}$ and $p_{j}$ to the point clouds. Then we define a reprojection loss to optimize the associated transformation as follows:
$L_{reproj}(i,j)=\|K\cdot(T_{ji}\cdot K^{-1}\cdot p_{i})-p_{j}\|,$ (8)
where $T_{ji}$ is the associated transformation from $i$ to $j$. By minimizing the overall reprojection loss for all pairs, we can roughly obtain the initial camera’s intrinsics and extrinsics.
3D Gaussians. We initialize the 3D Gaussians by 3D track points. We extract a set of tracks $\mathcal{P}$, where each element $(P,\{p_{i}\}_{i=1}^{l})\in\mathcal{P}$ represents a 3D track point $P$ and its corresponding matching points $\{p_{i}\}_{i=1}^{l}$ associated with the training images. The 3D track point $P$ is initialized as the centroid of the transformed projections of $\{p_{i}\}$:
$P=\frac{1}{l}\sum_{i=1}^{l}K^{-1}p_{i}.$ (9)
Notably, we use track points solely to initialize 3D Gaussians, as their positions will be refined by global optimization and constraints to accurately represent object surfaces.
4.2 Joint Optimization
4.2.1 Global Track Constraints
We leverage the global tracks obtained during initialization to enforce multi-view geometric consistency in joint optimization, both in 2D and 3D space. Our approach is intuitive and focuses on two key aspects. First, the reprojections of the 3D track points onto each image should closely match the original 2D feature points, ensuring that reprojection relationships are preserved throughout the optimization process. Second, the backprojection of the matched 2D feature points with 3DGS rendering depth from each training image should remain near their corresponding 3D track points in the scene, ensuring spatial consistency across all input images. Based on these two considerations, we introduce the 2D track loss and 3D track loss, respectively. Since the 3D track points in the scene should lie on the surface of the object, we address this by using the centroids of the initialized 3D Gaussians to represent the 3D track points. To achieve this, we introduce a scale loss to constrain the sizes of these Gaussians. This approach allows us to preserve the original function of the 3D Gaussians while treating them as virtual 3D spatial points. These points serve as the key elements for optimizing camera parameters and enhancing the global geometric consistency.
2D Track Loss. We reproject the 3D track points $P$ (Eq. 9) into the corresponding images using the associated camera parameters and compute the reprojection loss, which will be summed to calculate the total 2D track loss:
@@ 4.2.1 Global Track Constraints
3D Track Loss. We backproject the 2D feature points into 3D scene using the rendered depth and camera parameter associated with each point. The backprojection error is then computed with respect to the 3D tracked point $P$. Then errors are aggregated to calculate the overall 3D track loss:
@@
Scale Loss. In fact, 3D tracking points reside on object surfaces in the scene. By using 3D Gaussian centroids for tracking, it’s crucial to regulate their scale and align centroids closely with the actual object surfaces, ensuring accuracy in tracking point positions and minimizing projection errors. To achieve this, we incorporate a scale loss function to constrain the scale $S_{track}$ of these Gaussians $G_{track}$:
$L_{scale}=\sum_{S\in S_{track}}\|\max(S)\|.$ (12)
@@
$\mathcal{L}=$ $\lambda_{1}L_{1}+\lambda_{\mathrm{D-SSIM}}L_{\mathrm{D-SSIM}}$ (13)
$+\lambda_{t2d}L_{t2d}+\lambda_{t3d}L_{t3d}+\lambda_{scale}L_{scale}.$
4.2.2 Optimizing Camera Parameters
To optimize the camera parameters of 3D Gaussians simultaneously, the gradient of the loss function $\mathcal{L}$ respect to the camera parameters are needed. We derive these gradients accordingly, where the gradient of extrinsic parameters is:
@@ 4.2.2 Optimizing Camera Parameters
5 Experiments
5.1 Experimental Setup
Datasets. We conduct experiments on two real-world datasets, i.e., Tanks and Temples [12] and CO3D-V2 [20], and a Synthetic Dataset created by ourselves. Tanks and Temples, adapted from NoPe-NeRF [1], is used for novel view synthesis and pose estimation. It features 8 scenes with mild view changes, both indoors and outdoors. CO3D-V2 includes thousands of videos of various objects. Following the CF-3DGS [6], we select 8 scenes with significant camera movements to demonstrate our robustness. Synthetic Dataset comprises 4 scenes with about 150 frames each created using Blender [3], showcasing complex roaming and object-centric camera motions. It’s used to assess camera parameter estimation, providing ground truth for intrinsic and extrinsic parameters. For additional details on the synthetic dataset, see the supplementary material.
Note that, the Tanks and Temples dataset is a relatively simple test case for this task, as the camera motion in this scene is mostly linear or involves small movements, resulting in a very limited solution space for pose estimation. In contrast, the CO3D-V2 and our synthetic datasets contain much more complex camera trajectories. Therefore, as follows, we will emphasis more on the results of such challenging scenes, as they pose a higher level of difficulty.
Metrics. We use standard evaluation metrics, including PSNR, SSIM [28], and LPIPS [34] to evaluate the quality of novel view synthesis. For pose estimation, we rely on the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) [14, 1, 6]. $\mathrm{RPE}_{r}$ and $\mathrm{RPE}_{t}$ are utilized to measure the accuracy of rotation and translation, respectively. To ensure the metrics are comparable on the same scale, we align the camera poses using Umeyama’s method [26] for both estimation and evaluation. For camera focal length, we convert it to the field of view (FoV) and calculate the angular error, following [36].
Implementation Details. Our implementation is primarily based on gsplat [31], an accelerated 3DGS library. We implement joint optimization by backpropagating the gradient of camera parameters through modifications to the CUDA operator in the library. All parameters are optimized using Adam optimizer. For initialization, we optimize the relative pose between frames, and focal length. Then, the parameters of 3DGS, the absolute poses of cameras, and focal length are optimized. The camera pose is represented as a combination of an axis-angle representation $\mathrm{q}\in\mathfrak{so}(3)$ and a translation vector $\mathrm{t}\in\mathbb{R}^{3}$. All experiments are conducted on a single RTX 4090 GPU with 24GB VRAM.
Figure 3: Qualitative comparison for novel view synthesis on Tanks and Temples. We achieve better rendering results on details.
Figure 4: Qualitative comparison for novel view synthesis on CO3D-V2. We achieve the best rendering results on details.
Scenes Ours HT-3DGS CF-3DGS NoPe-NeRF
@@ 5.1 Experimental Setup
Tanks and Temples Church 29.39 0.92 0.10 31.34 0.94 0.08 30.23 0.93 0.11 25.17 0.73 0.39
Barn 31.98 0.94 0.08 34.95 0.97 0.05 31.23 0.90 0.10 26.35 0.69 0.44
Museum 31.92 0.94 0.08 31.59 0.95 0.08 29.91 0.91 0.11 26.77 0.76 0.35
Family 32.22 0.95 0.09 34.17 0.97 0.05 31.27 0.94 0.07 26.01 0.74 0.41
Horse 30.33 0.94 0.09 35.82 0.98 0.03 33.94 0.96 0.05 27.64 0.84 0.26
Ballroom 35.03 0.97 0.03 34.12 0.97 0.04 32.47 0.96 0.07 25.33 0.72 0.38
Francis 33.39 0.92 0.15 34.09 0.93 0.13 32.72 0.91 0.14 29.48 0.80 0.38
Ignatius 29.25 0.90 0.11 31.64 0.95 0.06 28.43 0.90 0.09 23.96 0.61 0.47
mean 31.68 0.94 0.09 33.53 0.96 0.07 31.28 0.93 0.09 26.34 0.74 0.39
CO3D-V2 34_1403_4393 28.68 0.88 0.21 32.52 0.93 0.14 27.75 0.86 0.20 28.62 0.80 0.35
46_2587_7531 31.83 0.92 0.12 30.65 0.91 0.13 25.44 0.80 0.21 25.30 0.73 0.46
106_12648_23157 26.18 0.83 0.19 23.43 0.73 0.28 22.14 0.64 0.34 20.41 0.46 0.58
110_13051_23361 33.44 0.94 0.11 29.95 0.87 0.19 29.69 0.89 0.29 26.86 0.73 0.47
245_26182_52130 33.82 0.93 0.20 28.59 0.87 0.27 27.24 0.85 0.30 25.05 0.80 0.49
407_54965_106262 28.73 0.86 0.39 - - - 27.80 0.84 0.35 25.53 0.83 0.58
415_57112_110099 30.37 0.88 0.22 27.23 0.78 0.30 22.14 0.64 0.34 20.41 0.46 0.58
429_60388_117059 25.70 0.70 0.35 - - - 24.44 0.68 0.36 22.19 0.62 0.56
mean 29.84 0.87 0.22 28.73 0.85 0.22 25.83 0.78 0.30 24.30 0.68 0.51

### v3

@@ Abstract
We present TrackGS, a novel method to integrate global feature tracks with 3D Gaussian Splatting (3DGS) for COLMAP-free novel view synthesis. While 3DGS delivers impressive rendering quality, its reliance on accurate precomputed camera parameters remains a significant limitation. Existing COLMAP-free approaches depend on local constraints that fail in complex scenarios. Our key innovation lies in leveraging feature tracks to establish global geometric constraints, enabling simultaneous optimization of camera parameters and 3D Gaussians. Specifically, we: (1) introduce track-constrained Gaussians that serve as geometric anchors, (2) propose novel 2D and 3D track losses to enforce multi-view consistency, and (3) derive differentiable formulations for camera intrinsics optimization. Extensive experiments on challenging real-world and synthetic datasets demonstrate state-of-the-art performance, with much lower pose error than previous methods while maintaining superior rendering quality. Our approach eliminates the need for COLMAP preprocessing, making 3DGS more accessible for practical applications.
Introduction
@@
Given a collection of images from a 3D scene along with the corresponding camera intrinsic and extrinsic parameters, 3D Gaussian Splatting (3DGS) (3DGS2023) can effectively represent the scene with a series of 3D Gaussians, and generate high-quality images from novel viewpoints. Due to its efficiency in training and superior performance in testing, 3DGS has become popular for a variety of applications including reconstruction, editing, and AR/VR etc. However, the effectiveness of 3DGS training relies on accurately pre-determined camera poses (i.e., camera extrinsics) and camera focal lengths (i.e., camera intrinsics). These parameters are typically derived using COLMAP (colmap2016) in advance. This preprocessing step is not only time-consuming but also impacts the training performance of 3DGS, particularly when dealing with complex camera movements and scenes.
Recent COLMAP-Free approaches (iNerf2021; SCNeRF2021; bian2022nopenerf; tracknerf2024eccv; fan2024instantsplat; CF-3DGS-2024; ji2024sfmfree3dgaussiansplatting) have tried to eliminate COLMAP dependencies by adding local constraints, such as photometric consistency or sequential frame alignment. While these methods work well for simple scenes with smooth camera trajectories, they struggle with more challenging cases—such as wide-baseline views, rapid camera movements, or unordered image collections—due to their reliance on incremental optimization and local geometric cues. This often leads to pose drift, scale ambiguity, and degraded rendering quality.
To address this problem, we propose TrackGS, a novel framework that integrates global track information into 3DGS to enforce multi-view geometric consistency robustly. Our key insight is that feature tracks provide strong, long-range constraints across disparate views, enabling stable optimization of both camera parameters and 3D Gaussians even in complex scenarios. Specifically, we introduce: 1. Track Gaussians, where a subset of Gaussians are anchored to 3D track points, ensuring their spatial accuracy through reprojection and backprojection losses. 2. A fully differentiable pipeline, that jointly optimizes camera intrinsics, extrinsics, and 3DGS parameters. We derive gradients for intrinsic parameters (e.g., focal length), enabling end-to-end training without any precomputed camera priors. Experimental results on both public and synthetic datasets demonstrate that our joint optimization framework with global track constraints outperforms previous methods.
@@
  • We propose a method to integrate track information with 3DGS, using global geometric constraints to simultaneously optimize camera parameters and 3DGS. To achieve this, we introduce 2D and 3D track losses to constrain reprojection and backprojection errors.
  • We propose a joint optimization framework for all camera parameters and 3D Gaussians. Without relying on any known camera parameters, we achieve full differentiability for the entire pipeline, seamlessly integrating camera parameters estimation, including both intrinsics and extrinsics, with 3DGS training.
@@
Figure 1: Comparisons on novel view synthesis and camera poses. We propose a novel 3DGS model without any known camera parameters by leveraging global track information. Compared with state-of-the-art methods, we provide higher rendering quality in novel view synthesis, and more accurate estimation of camera poses on benchmark datasets, including the challenging real-world indoor and outdoor or synthetic scenes with complicated camera movements (the right column).
Related Work
Novel View Synthesis
@@
Novel view synthesis is a foundational task in the computer vision and graphics, which aims to generate unseen views of a scene from a given set of images. Numerous methods have been developed to address this problem by approaching it as 3D geometry-based rendering, such as using meshes (worldsheet; FVS; SVS), MPI (MINE; single_view_mpi; stereo_magnification), point clouds (point_differentiable; point_nfs), etc.
@@
Figure 2: Overview. Given a set of images, our method obtains both camera intrinsics and extrinsics, as well as a 3DGS model. During the initialization, we extract the global tracks, and initialize camera parameters and Gaussians from image correspondences and monodepth. We determine Gaussian kernels with 3D track points, and then jointly optimize the parameters $K$, $T_{cw}$, 3DGS through the proposed global track constraints ($L_{t2d}$, $L_{t3d}$) and original photometric losses ($L_{1}$, $L_{D-SSIM}$).
@@
Recently, Neural Radiance Fields (NeRF) (2020NeRF) provide a novel solution to this problem by representing scenes as implicit radiance fields using neural networks, achieving photo-realistic rendering quality. Although having some works in improving efficiency (instant_nerf2022; lin2022enerf), the time-consuming training and rendering still limit its practicality. Alternatively, 3D Gaussian Splatting (3DGS) (3DGS2023) models the scene as explicit Gaussian kernels, with differentiable splatting for rendering. Its improved real-time rendering performance, lower storage and efficiency, quickly attract more attentions.
Optimizing Camera Poses in NeRFs and 3DGS
@@
Although NeRF and 3DGS can provide impressive scene representation, these methods all need accurate camera parameters (both intrinsic and extrinsic) as additional inputs, which are mostly obtained by COLMAP (colmap2016). When the prior is inaccurate or unknown, accurately estimating camera parameters and scene representations becomes crucial.
@@
In earlier studies, scene training and camera pose estimation relied solely on photometric constraints. iNeRF (iNerf2021) refines the camera poses using a pre-trained NeRF model. NeRFmm (wang2021nerfmm) introduces a joint optimization approach that simultaneously estimates camera poses and trains the NeRF model. BARF (barf2021) and GARF (2022GARF) propose a new positional encoding strategy to address the gradient inconsistency issues in positional embedding, achieving promising results. However, these methods only yield satisfactory optimization when the initial pose is very close to the ground truth, as photometric constraints alone can only enhance camera estimation quality within a limited range. Subsequently, SC-NeRF (SCNeRF2021) minimizes a projected ray distance loss based on correspondence between adjacent frames. NoPe-NeRF (bian2022nopenerf) utilizes monocular depth maps as geometric priors and defines undistorted depth loss and relative pose constraints.
@@
Regarding 3D Gaussian Splatting, CF-3DGS (CF-3DGS-2024) utilizes mono-depth information to refine the optimization of local 3DGS for relative pose estimation and subsequently learns a global 3DGS in a sequential manner. InstantSplat (fan2024instantsplat) targets sparse view scenes, initially employing MASt3R (mast3r_eccv24) to create a dense, pixel-aligned point set for initializing 3D Gaussian models and implements a parallel grid partitioning strategy to accelerate joint optimization. Jiang et al. (Jiang_2024sig) develops an incremental method for reconstructing camera poses and scenes, but struggles with complex scenes and unordered images. SFGS (ji2024sfmfree3dgaussiansplatting) interpolates frames for training and splits the scene into local clips, using a hierarchical strategy to build 3DGS model. It works well for simple scenes, but fails with dramatic motions due to unstable interpolation and low efficiency.
@@
where $c_{i}$ and $\alpha_{i}$ are the color and opacity of the Gaussians. Similarly, the depth of the scene perceived of a pixel is,
@@
Method
Overview. Given a set of images $\mathcal{I}=\{I_{i}\}_{i=1}^{M}$, with unknown extrinsic matrix $T_{cw,i}$ at each view and unknown intrinsic matrix denoted by $K$, our method aims to build a 3D Gaussian Splatting (3DGS) model while simultaneously estimating both the extrinsic and intrinsic matrices, as shown in Fig. 2. To achieve this goal, our key approach is to leverage the global track constraint to explicitly capture and enforce multi-view geometric consistency, which serves as the foundation for accurately estimating both the 3DGS model and the camera parameters. During initialization, we extract 2D feature correspondences and derive global tracks using off-the-shelf methods. We then initialize both the camera parameters and the original 3D Gaussians with the estimated 3D track points. Subsequently, we extract a subset of Gaussian kernels from the vanilla model, referred to as track Gaussians, to integrate the track-based geometric prior into 3DGS. Building on this, we further propose a joint optimization framework with two additional loss terms: a 2D track loss and a 3D track loss. The 2D track loss enforces multi-view geometric consistency by minimizing the error between the reprojected track pixels and their references. The 3D track loss constrains the track Gaussians to remain aligned with the scene surface by penalizing the distance between the back-projected 3D points (from rendered depth) and the track Gaussians. We derive and implement the differentiable components of the camera parameters, including both the extrinsic and intrinsic matrices. This allows us to apply the chain rule, enabling seamless joint optimization of the 3DGS model and the camera parameters.
Initialization
Since the training set contains only images, we need to extract essential information during initialization. To initialize the 3D Gaussians, we estimate monocular depth maps $D$ for each image $I$ using depth estimators such as DPT (dpt2021iccv), following the setups in CF-3DGS (CF-3DGS-2024) and SFGS (ji2024sfmfree3dgaussiansplatting). In addition, to construct global track constraints for later optimization, we detect 2D feature points $\{p_{i}\}$ in each image $I$ and compute feature correspondences across all images using off-the-shelf methods (detone2018superpoint; sarlin20superglue).
@@
Given the estimated mono-depth maps $D_{i},D_{j}$ and 2D correspondences $p_{i},p_{j}$, the relative transformation $T_{ij}$ between frames $i$ and $j$ is computed by aligning point clouds $p_{i}^{*},p_{j}^{*}$, where $p^{*}=D(p)\cdot K^{-1}\cdot p$. These alignments provide a coarse initialization of both camera intrinsics and extrinsics.
Global Tracks. By using Union-Find algorithm over feature points, We extract a set of tracks $\mathcal{P}$, where each element $(P,\{p_{i}\}_{i=1}^{l})\in\mathcal{P}$ represents a 3D track point $P$ and its corresponding matching points $\{p_{i}\}_{i=1}^{l}$ associated with the training images. With the estimated camera parameters and mono-depth maps, the 3D point $P$ is initialized as the average position of $\{p_{i}^{*}\}_{i}$. Notably, we use track points solely to initialize 3D Gaussians, as their positions will be refined by global optimization and constraints to accurately represent object surfaces.
Joint Optimization
Track Gaussians and Global Track Constraints.
During joint optimization, we leverage the global tracks obtained during initialization to enforce multi-view geometric consistency in both 2D and 3D space. Our method is built on two key components. First, to associate the tracks with the 3DGS model, we select a subset of 3D Gaussians associated with each track, referred to as track Gaussians, whose centroids coincide with the 3D track points and serve as geometric anchors within the model. Second, we introduce two additional loss terms to impose global track constraints over these track Gaussians. The reprojection loss (2D track loss) enforces that the reprojections of 3D track points on each image closely aligned with the corresponding 2D feature points. The backprojection loss (3D track loss) enforces that the matched 2D feature points, when back-projected using the 3DGS rendering depth, remain close to the same corresponding 3D track points across different views. Beyond enforcing multi-view consistency, the backprojection loss further ensures that the track Gaussians treated as geometric anchors, as the projected feature points with rendering depth are considered to lie on the scene surface. These points serve as key elements for optimizing camera parameters and enhancing the global geometric consistency.
2D Track Loss. We reproject the 3D track point $P$ into the corresponding images using the associated camera parameters and compute the reprojection loss, which will be summed to calculate the total 2D track loss:
@@
3D Track Loss. We backproject the 2D feature points into 3D scene using the rendered depth and camera parameters associated with each point. The backprojection error is then computed with respect to the 3D track point $P$. Then errors are aggregated to calculate the overall 3D track loss:
@@
$\mathcal{L}=(1-\lambda)L_{1}+\lambda L_{\mathrm{D-SSIM}}+\lambda_{t2d}L_{t2d}+\lambda_{t3d}L_{t3d}.$ (10)
Optimizing Camera Parameters.
To optimize the camera parameters of 3D Gaussians simultaneously, the gradient of the loss function $\mathcal{L}$ with respect to the camera parameters are needed. We derive these gradients accordingly, where the gradient of extrinsic parameters is:
@@
Figure 3: Qualitative comparison for NVS on Tanks and Temples. We achieve better rendering results on details.
T&T Church Horse Ignatius Ballroom
@@
CF-3DGS 21.96 0.71 0.27 17.36 0.64 0.33 14.96 0.31 0.59 22.20 0.72 0.25
SFGS 22.46 0.72 0.22 20.86 0.75 0.19 12.16 0.32 0.72 17.42 0.49 0.27
Ours 25.56 0.84 0.16 27.60 0.90 0.12 22.12 0.71 0.22 25.94 0.86 0.12
CO3D V2 46_2587_7531 110_13051_23361 245_26182_52130 415_57112_110099
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
CF-3DGS 25.44 0.80 0.21 29.69 0.89 0.29 27.24 0.85 0.30 22.14 0.64 0.34
SFGS 30.65 0.91 0.13 29.95 0.87 0.19 28.59 0.87 0.27 27.23 0.78 0.30
Ours 31.83 0.92 0.12 33.44 0.94 0.11 33.82 0.93 0.20 30.37 0.88 0.22
Synthetic classroom lego_c2 livingroom bedroom
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
CF-3DGS 19.69 0.69 0.46 15.93 0.31 0.55 16.63 0.57 0.57 16.98 0.65 0.45
SFGS 16.79 0.67 0.55 13.64 0.29 0.68 15.11 0.49 0.60 12.39 0.54 0.54
Ours 36.26 0.94 0.15 29.36 0.90 0.12 33.52 0.88 0.24 31.17 0.93 0.13